# Improving Medical Visual Representation Learning with Pathological-level Cross-Modal Alignment and Correlation Exploration

### v1

@@ 4.3 Results
Results on Classification and Image-to-text Retrieval. Results in Table 1 demonstrate the effective of PLACE on image classification tasks where our model achieves the best performance on both the fine-tuned settings and zero-shot scenario, outperforming the second-best results by a significant margin. Moreover, PLACE remarkably improves the zero-shot classification performance on both the CheXpert and COVIDx dataset. A similar pattern is shown in the image-to-text retrieval task in Table 3 where our model obtains the highest scores for all the setting of $K$, indicating better capability of aligning the pathology information. These results, especially for the zero-shot scenarios, further confirm the efficacy of PLACE.
@@
Results on Semantic Segmentation. We report the results of semantic segmentation in Table 1. As can be seen, PLACE surpasses the previous methods by a notable margin on both the RSNA and SIIM datasets. Notably, the superiority of PLACE becomes more pronounced in scenarios characterized by limited data availability, e.g., 1% and 10% of the training samples, suggesting the efficacy of our approach in learning highly generalizable, fine-grained representations, which is particularly crucial in small-data regimes on tasks with higher demand for localized and fine-grained features.
@@
Results on Report Generation. We compares the efficacy of PLACE against recent RRG-specific models (including VLP specific to MRG), and general Visual Language Pretraining (VLP) approaches. Furthermore, to underscore the efficacy of PLACE, we present results of the same models with visual extractor/encoder weights initialized from ImageNet pre-trained models. As illustrated in Table 5, our approach significantly outperforms previous VLP methods and demonstrates superior performance compared to models initialized with ImageNet weights. These findings highlight the cross-modal capabilities of PLACE and its effectiveness in learning a fine-grained and pathology-enriched visual representation. Moreover, even when utilizing a frozen weight setting, our model achieves competitive results compared to most models tailored to MRG tasks. It is noteworthy that our evaluation of the effectiveness of PLACE is based on standard transformer-based baselines, without incorporating advanced MRG-specific techniques or additional knowledge sources. Despite these achievements, there remains a performance disparity between general VLP models and those designed specifically for MRG tasks. This could be attributed to two main factors. Firstly, the data volume during the pre-training phase is comparable to that of the MRG task since both utilize the MIMIC-CXR dataset, leading to limited additional knowledge transfer when adapting general VLP models to MRG tasks. Secondly, MRG-specific methods often leverage supplementary information or data sources such as disease labels and knowledge graphs. For example, ATL-CAC incorporates crucial clinical history, such as comparison and indication, as auxiliary inputs to enhance model performance.
@@ 4.4.1 Contribution of each component
We investigate the contribution of each proposed module, including the Pathological-level Cross-modal Alignment (PCMA) and Cross-modal Correlation Exploration (CCE) on semantic segmentation, object detection, and image classification tasks. The model optimized only by instance-level cross-modal alignment (ICMA) is considered as the baseline. As shown in Table 4, each stage of our design shows consistent improvements over the baseline model. Additionally, further combining the PCMA and CCE modules brings more remarkable gains for downstream tasks, especially for challenging ones such as semantic segmentation requiring fine-grained details and zero-shot classification. It is noteworthy that our method shows higher efficacy in the setting when fewer training samples available. For example, our full PLACE model enhances the performance of baseline by 4.6 (RSNA) and 10.1 (SIIM) on the segmentation task in the 1% training scenario, and significantly improves the accuracy of the zero-shot classification from 36.5% to 44.0%. These results show that PLACE can learn a more generalizable and fine-grained visual representation, confirming the effectiveness of our proposed methods. We attribute the improvements mainly to the better exploitation of the naturally exhibited pathological correspondences across image and reports, and to the exploration of the fine-grained details.
@@ 4.4.2 Additional exploration of the PCMA objective
In addition to the ablation study presented in the previous section, we conduct a more comprehensive investigation of the effectiveness of our proposed PCMA objective by comparing our complete PLACE with two variants: (1) PLACE without V-PORs: in the PCMA module, we remove the V-PORs, thus downgrading the alignment to the level between the original localized visual tokens and T-PORs; (2) PLACE without T-PORs: in the PCMA module, we remove the T-PORs, thus downgrading the alignment to the level between the original localized visual tokens and T-PORs. As evidenced in Table 6, the exclusion of V-PORs or T-PORs results in diminished performance, and a more significant decline can be observed in more complex dense prediction tasks, such as semantic segmentation. These findings corroborate the efficacy of our V-PORs design and the proposed alignment module (a.k.a. PCMA Align) on the pathological level. Moreover, they underscore the significance of augmenting the granularity from image patches $\leftrightarrow$ words to pathological regions $\leftrightarrow$ pathological sentences (a.k.a., textual pathological observations).
@@
We investigate the influence of PLACE to the number of pathology query tokens $N_{p}$ by varying $N_{p}$ from $10$ to $16$ on semantic segmentation and classification tasks. The results in Table 7 show that PLACE is relatively robust to $N_{p}$. Nevertheless, it is still beneficial to set an appropriate value to achieve the best performance, since $N_{p}$ represents the number of pathological observations that the model will extract from the images. A too large value exceeding the total pathological observations may introduce noisy information, while a too small a value struggles to cover all the pathological observations.
@@
We explore the influence of PLACE to the patch size $PS$ in the CCE module on semantic segmentation and classification tasks. Note that the image size needs to be wholly divisible by $PS$ to ensure that the total number of patches is an integer. We conduct the experiments on a $PS$ of $16,28,32,56$, resulting in a total of $196$,$64$,$49$ and $16$ image patches attending the calculation of covariance matrix. As can be seen in Table 8, PLACE is not overly sensitive to the value of the patch size. Nonetheless, a smaller value may disrupt the integrity of semantic information, potentially compromising the efficacy of the covariance matrix in serving as a correlation descriptor. Conversely, excessively large values may introduce non-discriminative noise. Hence, an appropriate value of $PS$ can bring more significant performance to PLACE.
$PS$ RSNA(Dice) COVIDx(ACC)
@@
Table 8: Effect of varying $PS$, the patch size in the Cross-Modal Correlation Exploration Module.
@@ 4.5 Computational Complexity Analysis
Here, we analyze the parameter count of the Base model and the proposed PLACE models as shown in Table 9. The base model comprises solely the image and text encoders. By incorporating the full PLACE module (Base+PCMA+CCE), an increase of 13.0M learnable parameters is observed. Despite the rise in computational complexity, it remains relatively insignificant compared to the base model (113.4M), while significantly enhancing model performance across various downstream tasks. Additionally, the integration of the CCE module introduces only 0.9M learnable parameters to the models, which is negligible in comparison to the base model, yet lead to notable enhancements in tasks such as medical detection and segmentation.
The CCE module is constructed with a linear layer sized $D\times K$, where $D$ represents the dimension of the global report representation (768 here) and $K$ signifies the number of elements in the covariance matrix. It is noteworthy that the covariance matrix assesses the correlation between each unique pair of patches. As a result, the value of $K$ increases quadratically with the number of image patches, which in turn is also quadratically to the patch size with a fixed image resolution. For example, given the image resolution of 224 and the patch size $P$, the total number of element in covariance matrix $K=(\lfloor\frac{224}{P}\rfloor)^{4}$. Consequently, the number of parameters in the CCE module grows inversely proportional to the fourth power of the patch size. Given that the covariance matrix is symmetric, it is only necessary for the model to predict the upper triangular portion of the matrix (correlation for image patch to itself is also excluded). This reduces the parameter count by half to $\{\frac{K}{2}-{num\_patches}\}$ in our implementation. Considering a small patch size of 16, this results in a total of 14.7M learnable parameters. However, as indicated in the ablation study in subsection 4.4 (4), the model is not overly sensitive to variations in patch size. It is generally advised against choosing very small patch sizes like 8 or 16 due to the risk of disrupting semantic continuity. Conversely, selecting a very large patch size, for instance, 112, may compromise the exploration of fine-grained local information. Therefore, it remains crucial to choose an appropriate patch size to obtain the optimal performance and balance better performance with higher efficiency. A commonly adopted configuration in various downstream applications, i.e., the one used in our study with a patch size of $32$ and an image resolution of $224\times 224$, introduces only 0.9M parameters.
@@ 5.0.1 Visual pathology observations
To further explore whether our proposed method, PLACE, can truly extract useful visual pathology observations, we present a T-SNE visualisation of visual pathology observation representations from 100 samples in Figure 3 where points of the same colour are retrieved from the same pathology query token. Obviously, V-PORs retrieved by the same pathology query token are clustered, indicating that each is specifically functioned to extract one type of pathology from the visual tokens as expected. It is worth noting that in addition to the large cluster for different pathogenesis, each pathological-level cluster contains several sub-clusters, aligning to real-world situations that each pathology may contain multiple patterns such as normal, abnormal, and those of different severities. To substantiate that these V-PORs can effectively extract information related to atelectasis from the images, we present visualizations of the attention scores between the V-POR that exhibits the highest similarity to the atelectasis-related T-PORs, and every local visual token produced by the cross-attention module in the VPOE shown in Figure 5. Evidently, these V-PORs successfully activate the regions associated with atelectasis in the image. Furthermore, they even identify precise locations as referenced in left lower lobe, bilateral lower lobe and bibasilar. The preceding results unequivocally demonstrate that each pathology query token possesses the ability to extract pathology observations from the images and exhibit pathological-level alignment through the reasonable and meticulous design of our VPOE and PCMA modules, thereby further confirming our theory.
@@ 5.0.2 V-PORs and Pathological-Level Alignment
    The query tokens facilitate the extraction of consistent V-PORs across various images &\&& The V-PORs concentrate on the same anatomical regions $\rightarrow$ The V-PORs demonstrate an ability to capture pathological observations pertinent to a specific anatomical location.
@@
    The query tokens effectively extract V-PORs that are most relevant to the T-PORs. Given that the T-PORs, learned directly from the text, are clinically significant, one may conclude that the proposed PCMA module functions effectively, i.e., A&\&&B $\rightarrow$ C.
The analysis in the previous paragraphs has confirmed the condition A. To validate B, we randomly select 200 samples showing the presence of the Atelectasis and locate the relevant sentences describing the Atelectasis observation by checking whether the sentence contains the pathology observation name [^1] Samples without comprising the observation name are simply removed.. Subsequently, we record the index of the V-POR that demonstrates the greatest similarity with each selected T-POR from the entire set of samples. Following this, we evaluate the consistency of the selected V-POR across all samples by calculating the proportion of the most frequently selected V-POR (top-1 result). The findings indicate that $47.4\%$ of the samples exhibit a consistent top-one-ranked V-POR. Acknowledging that certain sentences may encompass multiple observations, i.e., sentence “extremely low lung volumes with pulmonary vascular crowding and left basilar atelectasis,” we additionally present the top-2 result ($59.3\%$), which includes samples wherein the most frequently selected V-POR ranks second. The preceding results unequivocally demonstrate that the query tokens possess the ability to extract pathology observations from the images and exhibit pathological-level alignment through the reasonable and meticulous design of our PCMA module, thereby further confirming our framework.
@@ 5.0.4 Visualization of Important Words
We present the ten most significant words (highlighted in red) from four sample reports learned by the model in Table 10. These words are identified by averaging the attention weights from BERT’s final layer. As can be seen, the majority of these highlighted words, i.e., consolidation, pneumonia, pleural, minimal, pertain to patients’ medical conditions. Furthermore, our model demonstrates a capability of discerning the location and severity of pathological observations, as illustrated by words such as mild, stable, lower, left, right. These findings suggest that our proposed framework, PLACE, is capable of acquiring a more fine-grained and pathologically enriched representation.
@@
Table 10: An illustration of the top ten important words of four samples learnt by the text encoder of PLACE during the training process. The important words are highlighted in red colour.
@@ 6 Conclusions
This paper introduces PLACE, a cutting-edge framework that enhances cross-modal alignment at the level of pathological observations for learning medical visual representation from image-report joint-training. This is achieved by a novel cross-modal pathological-level alignment module without the need for additional human annotations. Additionally, we develop a new proxy task, termed Cross-modal Correlation Exploration, which enables our model to capture more critical fine-grained details. These carefully and reasonably designed modules within PLACE allow it to learn a more generalizable and fine-grained medical visual representation. Experimental evaluations across multiple downstream tasks demonstrate the superiority of PLACE over previous state-of-the-art models, offering a robust solution to automatically align pathological information between images and reports without reliance on external human annotations.
@@
Our objective is to enhance medical visual representation learning by improving pathological-level cross-modal alignment and fine-grained representation learning, allowing seamless adaptation to different datasets or domains without requiring additional annotations such as bounding boxes or disease labels. Therefore, an explicit mechanism for distinguishing the abnormalities in the observation alignment is not in place; instead, the reliance is on the proposed PCMA module and T-POR to implicitly navigate through pathological conditions. However, there might be inconsistencies in aligning pathological conditions due to imperfections in the solution, e.g., normal V-POR to abnormal T-POR, even though the PCMA module at least enforces consistency by ensuring that pathological observations related to a specific anatomical region align correctly within samples. Furthermore, the number of parameters in the CCE module increases inversely with the fourth power of the patch size, potentially restricting applications that necessitate large image resolution using a small patch size. Our future work will focus on addressing these limitations without requiring additional annotations.
Acknowledgment

### v2

@@ 4.3 Results
Results on Classification and Image-to-text Retrieval. Results in Table 1 demonstrate the effective of PLACE on image classification tasks where our model achieves the best performance on both the fine-tuned settings and zero-shot scenario, outperforming the second-best results by a significant margin. Moreover, PLACE improves remarkably the zero-shot classification performance on both the CheXpert and COVIDx dataset. A similar pattern is shown in the image-to-text retrieval task in Table 3 where our model obtains the highest scores for all the setting of $K$, indicating better capability of aligning the pathology information. These results, especially for the zero-shot scenarios, further confirm the efficacy of PLACE.
@@
Results on Semantic Segmentation. We report the results of semantic segmentation in Table 1. As can be seen, PLACE surpasses previous methods by a notable margin on both the RSNA and SIIM datasets. Notably, the superiority of PLACE becomes more pronounced in scenarios characterized by limited data availability, e.g., 1% and 10% of the training samples, suggesting the efficacy of our approach in learning highly generalizable, fine-grained representations, which is particularly crucial in small-data regimes on tasks with higher demand for localized and fine-grained features.
@@
Results on Report Generation. We compare the efficacy of PLACE against recent RRG-specific models (including VLP specific to MRG), and general Visual Language Pretraining (VLP) approaches. Furthermore, to underscore the efficacy of PLACE, we present results of the same models with visual extractor/encoder weights initialized from ImageNet pre-trained models. As illustrated in Table 5, our approach significantly outperforms previous VLP methods and demonstrates superior performance compared to models initialized with ImageNet weights. These findings highlight the cross-modal capabilities of PLACE and its effectiveness in learning a fine-grained and pathology-enriched visual representation. Moreover, even when utilizing a frozen weight setting, our model achieves competitive results compared to most models tailored to MRG tasks. It is noteworthy that our evaluation of the effectiveness of PLACE is based on standard transformer-based baselines, without incorporating advanced MRG-specific techniques or additional knowledge sources. Despite these achievements, there remains a performance disparity between general VLP models and those designed specifically for MRG tasks. This could be attributed to two main factors. Firstly, the data volume during the pre-training phase is comparable to that of the MRG task since both utilize the MIMIC-CXR dataset, leading to limited additional knowledge transfer when adapting general VLP models to MRG tasks. Secondly, MRG-specific methods often leverage supplementary information or data sources such as disease labels and knowledge graphs. For example, ATL-CAC incorporates crucial clinical history, such as comparison and indication, as auxiliary inputs to enhance model performance.
@@ 4.4.1 Contribution of each component
We investigate the contribution of each proposed module, including the Pathological-level Cross-modal Alignment (PCMA) and Cross-modal Correlation Exploration (CCE) on semantic segmentation, object detection, and image classification tasks. The model optimized only by instance-level cross-modal alignment (ICMA) is considered as the baseline. As shown in Table 4, each stage of our design shows consistent improvements over the baseline model. Additionally, further combining the PCMA and CCE modules brings more notable gains for downstream tasks, especially for challenging ones such as semantic segmentation requiring fine-grained details and zero-shot classification. It is noteworthy that our method shows higher efficacy in the setting when fewer training samples available. For example, our full PLACE model enhances the performance of baseline by 4.6 (RSNA) and 10.1 (SIIM) on the segmentation task in the 1% training scenario, and significantly improves the accuracy of the zero-shot classification from 36.5% to 44.0%. These results show that PLACE can learn a more generalizable and fine-grained visual representation, confirming the effectiveness of our proposed methods. We attribute the improvements mainly to the better exploitation of the naturally exhibited pathological correspondences across image and reports, and to the exploration of the fine-grained details.
@@ 4.4.2 Additional exploration of the PCMA objective
In addition to the ablation study presented in the previous section, we conduct a more comprehensive investigation of the effectiveness of our proposed PCMA objective by comparing our complete PLACE with two variants: (1) PLACE without V-PORs: in the PCMA module, we remove the V-PORs, thus downgrading the alignment to the level between the original localized visual tokens and T-PORs; (2) PLACE without T-PORs: in the PCMA module, we remove the T-PORs, thus downgrading the alignment to the level between the original localized visual tokens and T-PORs. As evidenced in Table 6, the exclusion of V-PORs or T-PORs results in diminished performance, and a more significant decline can be observed in more complex dense prediction tasks, such as semantic segmentation. These findings corroborate the efficacy of our V-PORs design and the proposed alignment module (namely PCMA Align) on the pathological level. Moreover, they underscore the significance of augmenting the granularity from image patches $\leftrightarrow$ words to pathological regions $\leftrightarrow$ pathological sentences (the textual pathological observations).
@@
We investigate the influence of PLACE to the number of pathology query tokens $N_{p}$ by varying $N_{p}$ from $10$ to $16$ on semantic segmentation and classification tasks. The results in Table 7 show that PLACE is relatively robust to $N_{p}$. Nevertheless, it is still beneficial to set an appropriate value to achieve the best performance, since $N_{p}$ represents the number of pathological observations that the model will extract from the images. A too large a value exceeding the total pathological observations may introduce noisy information, while a too small a value struggles to cover all the pathological observations.
@@
We explore the influence on PLACE of the patch size $PS$ in the CCE module, on semantic segmentation and classification tasks. Note that the image size needs to be wholly divisible by $PS$ to ensure that the total number of patches is an integer. We conduct the experiments on a $PS$ of $16,28,32,56$, resulting in a total of 196196,6464,$49$ and $16$ image patches attending the calculation of covariance matrix. As can be seen in Table 8, PLACE is not overly sensitive to the value of the patch size. Nonetheless, a smaller value may disrupt the integrity of semantic information, potentially compromising the efficacy of the covariance matrix in serving as a correlation descriptor. Conversely, excessively large values may introduce non-discriminative noise. Hence, an appropriate value of $PS$ can bring more significant performance to PLACE.
PSPS RSNA(Dice) COVIDx(ACC)
@@
Table 8: Effect of varying PSPS, the patch size in the Cross-Modal Correlation Exploration Module.
@@ 4.5 Computational Complexity Analysis
Here, we analyze the parameter count of the Base model and the proposed PLACE models as shown in Table 9. The base model comprises solely the image and text encoders. By incorporating the full PLACE module (Base+PCMA+CCE), an increase of 13.0M learnable parameters is observed. Despite the rise in computational complexity, it remains relatively insignificant compared to the base model (113.4M), while significantly enhancing model performance across various downstream tasks. Additionally, the integration of the CCE module introduces only 0.9M learnable parameters to the models, which is negligible in comparison to the base model, yet leads to notable enhancements in tasks such as medical detection and segmentation.
The CCE module is constructed with a linear layer sized $D\times K$, where $D$ represents the dimension of the global report representation (768 here) and $K$ signifies the number of elements in the covariance matrix. It is noteworthy that the covariance matrix assesses the correlation between each unique pair of patches. As a result, the value of $K$ increases quadratically with the number of image patches, which in turn is also quadratic to the patch size with a fixed image resolution. For example, given the image resolution of 224 and the patch size $P$, the total number of element in covariance matrix $K=(\lfloor\frac{224}{P}\rfloor)^{4}$. Consequently, the number of parameters in the CCE module grows inversely proportional to the fourth power of the patch size. Given that the covariance matrix is symmetric, it is only necessary for the model to predict the upper triangular portion of the matrix (correlation for image patch to itself is also excluded). This reduces the parameter count by half to $\{\frac{K}{2}-{num\_patches}\}$ in our implementation. Considering a small patch size of 16, this results in a total of 14.7M learnable parameters. However, as indicated in the ablation study in subsection 4.4 (4), the model is not overly sensitive to variations in patch size. It is generally advised against choosing very small patch sizes like 8 or 16 due to the risk of disrupting semantic continuity. Conversely, selecting a very large patch size, for instance, 112, may compromise the exploration of fine-grained local information. Therefore, it remains crucial to choose an appropriate patch size to obtain the optimal performance and balance better performance with higher efficiency. A commonly adopted configuration in various downstream applications, i.e., the one used in our study, is with a patch size of $32$ and an image resolution of $224\times 224$, introduces only 0.9M parameters.
@@ 5.0.1 Visual pathology observations
To further explore whether PLACE can truly extract useful visual pathology observations, we present a T-SNE visualisation of visual pathology observation representations from 100 samples in Figure 3 where points of the same colour are retrieved from the same pathology query token. Obviously, V-PORs retrieved by the same pathology query token are clustered, indicating that each is specifically functioned to extract one type of pathology from the visual tokens as expected. It is worth noting that in addition to the large clusters for different pathogenesis, each pathological-level cluster contains several sub-clusters, aligning to real-world situations that each pathology may contain multiple patterns such as normal, abnormal, and those of different severities. To substantiate that these V-PORs can effectively extract information related to atelectasis from the images, we present visualizations of the attention scores between the V-POR that exhibits the highest similarity to the atelectasis-related T-PORs, and every local visual token produced by the cross-attention module in the VPOE shown in Figure 5. Evidently, these V-PORs successfully activate the regions associated with atelectasis in the image. Furthermore, they even identify precise locations as referenced in left lower lobe, bilateral lower lobe and bibasilar. The preceding results unequivocally demonstrate that each pathology query token possesses the ability to extract pathology observations from the images and exhibit pathological-level alignment through the meticulous design of our VPOE and PCMA modules, thereby further confirming our theory.
@@ 5.0.2 V-PORs and Pathological-Level Alignment
    The query tokens facilitate the extraction of consistent V-PORs across various images &\& The V-PORs concentrate on the same anatomical regions $\rightarrow$ The V-PORs demonstrate an ability to capture pathological observations pertinent to a specific anatomical location.
@@
    The query tokens effectively extract V-PORs that are most relevant to the T-PORs. Given that the T-PORs, learned directly from the text, are clinically significant, one may conclude that the proposed PCMA module functions effectively, i.e., A&\&B $\rightarrow$ C.
The analysis in the previous paragraphs has confirmed the condition A. To validate B, we randomly select 200 samples showing the presence of the Atelectasis and locate the relevant sentences describing the Atelectasis observation by checking whether the sentence contains the pathology observation name [^1] Samples without comprising the observation name are simply removed.. Subsequently, we record the index of the V-POR that demonstrates the greatest similarity with each selected T-POR from the entire set of samples. Following this, we evaluate the consistency of the selected V-POR across all samples by calculating the proportion of the most frequently selected V-POR (top-1 result). The findings indicate that $47.4\%$ of the samples exhibit a consistent top-one-ranked V-POR. Acknowledging that certain sentences may encompass multiple observations, e.g., sentence “extremely low lung volumes with pulmonary vascular crowding and left basilar atelectasis,” we additionally present the top-2 result ($59.3\%$), which includes samples wherein the most frequently selected V-POR ranks second. The preceding results unequivocally demonstrate that the query tokens possess the ability to extract pathology observations from the images and exhibit pathological-level alignment with our PCMA module, thereby further substantiating the validity of our framework.
@@ 5.0.4 Visualization of Important Words
We present the ten most significant words (highlighted in bold) from four sample reports learned by the model in Table 10. These words are identified by averaging the attention weights from BERT’s final layer. As can be seen, the majority of these highlighted words, i.e., consolidation, pneumonia, pleural, minimal, pertain to patients’ medical conditions. Furthermore, our model demonstrates a capability of discerning the location and severity of pathological observations, as illustrated by words such as mild, stable, lower, left, right. These findings suggest that our proposed framework, PLACE, is capable of acquiring a more fine-grained and pathologically enriched representation.
@@
Table 10: An illustration of the top ten important words of four samples learnt by the text encoder of PLACE during the training process. The important words are highlighted in bold.
@@ 6 Conclusions
This paper introduces PLACE, a cutting-edge framework that enhances cross-modal alignment at the level of pathological observations for learning medical visual representation from image-report joint-training. This is achieved by a novel cross-modal pathological-level alignment module without the need for additional human annotations. Additionally, we develop a new proxy task, termed Cross-modal Correlation Exploration, which enables our model to capture more critical fine-grained details. These reasonable and carefully designed modules within PLACE allow it to learn a more generalizable and fine-grained medical visual representation. Experimental evaluations across multiple downstream tasks demonstrate the superiority of PLACE over previous state-of-the-art models, offering a robust solution to automatically align pathological information between images and reports without reliance on external human annotations.
@@
Our objective is to enhance medical visual representation learning by improving pathological-level cross-modal alignment and fine-grained representation learning, allowing seamless adaptation to different datasets or domains without requiring additional annotations such as bounding boxes or disease labels. Therefore, an explicit mechanism for distinguishing the abnormalities in the observation alignment is not present; instead, the reliance is on the proposed PCMA module and T-POR to implicitly navigate through pathological conditions. Thus there might be inconsistencies in aligning pathological conditions due to imperfections in the solution, e.g., normal V-POR to abnormal T-POR, even though the PCMA module at least enforces consistency by ensuring that pathological observations related to a specific anatomical region align correctly within samples. Furthermore, the number of parameters in the CCE module increases inversely with the fourth power of the patch size, potentially restricting applications that necessitate large image resolution using a small patch size. Our future work will focus on addressing such limitations without requiring additional annotations.
Acknowledgments